What is the sum of the first 10 numbers of the arithmetic sequence 1,2,3,……
 
Sum of how many terms of this sequence starting from the first is 300?​

Grade:12

1 Answers

Vikas TU
3 years ago
Dear student 
a=1
d=1
n=10 
 S10=10/2(2*1+(10-1)1)|
  S10=5(2+9)
 S10=5*11
 S10=55
When n=300
S300=300/2(2*1+(300-1)1)
S300=150(2+299).
S300=150*311
S300=46650